# Numerically “exact” charge transport dynamics in a dissipative electron–phonon model rationalizing the success of the transient localization scenario

### v1

@@ Abstract
Optical conductivity in molecular semiconductors is suppressed in the terahertz region, featuring the displaced Drude peak that reflects carriers’ transient localization (TL) by slow intermolecular vibrations. Meanwhile, recent computations in minimal models evidence optical-conductivity enhancements below the characteristic vibrational frequency, which cannot be captured by the TL phenomenology. These models assume undamped vibrations, whereas atomistic simulations show that the vibrations most strongly modulating carrier’s hopping amplitude are damped. Here, we use the dissipaton equations of motion (DEOM) method to compute finite-temperature real-time current autocorrelation function in a one-dimensional model with Brownian-oscillator spectral density of nonlocal carrier–phonon interaction. We exploit the dissipaton algebra to handle the phonon-assisted current, reduce the method’s computational requirements by working in momentum space, and confirm that numerically stable transport dynamics are virtually independent of a specific DEOM closing scheme. With increasing damping, we find that DEOM optical-conductivity profiles become increasingly qualitatively similar to TL predictions. For parameters representative of room-temperature hole transport in single-crystal rubrene, we conclude that the TL phenomenology is established already in the underdamped-oscillator regime. Reasonable variations in the damping constant weakly affect the carrier mobility, which remains within experimental bounds. Overall, our results strongly suggest that optical-conductivity enhancements at very low frequencies are artifacts of undamped vibrations, and rationalize the success of the TL phenomenology in describing experimental data.
@@ I Introduction
This state of affairs has changed very recently with us devising a numerically “exact” hierarchical equations of motion (HEOM)-based approach to carrier transport [24, 25] in the above-described one-dimensional Peierls (or Su–Schrieffer–Heeger) [26, 27, 28] model. For model parameters appropriate for room-temperature hole transport along the maximum-conductivity direction in single crystals of rubrene, [23] we evidence subtle qualitative differences between HEOM results and TLS predictions at long times (low frequencies). Namely, HEOM (TLS) shows that the diffusive transport is ultimately reached from the super-diffusive (sub-diffusive) side, with optical response increasing (decreasing) as $\omega\to 0$[25] The results of the most recent quantum–classical approaches, [21] to which the HEOM method reduces in the slow-phonon limit, display long-time and low-frequency features that are fully compatible with our HEOM results. Meanwhile, most experimental low-frequency signatures align with TLS (Drude–Anderson) predictions, [9, 10, 11, 12] though there are exceptions. [8]
@@
The aforementioned minimal-model studies [25, 21] assume that the carrier hopping amplitude is modulated by a single phonon mode (also known as the “killer” mode [7, 29]) of frequency $\omega_{0}\approx 50\>\mathrm{cm}^{-1}$. Although the “killer” mode does have the dominant contribution to the total dynamic disorder, the remaining low-frequency vibrations, typically quasi-continuously distributed over the $(0-200)\>\mathrm{cm}^{-1}$ range, [30] act as a thermal bath that effectively introduces friction to the “killer”-mode dynamics. [30, 31] Indeed, molecular-dynamics simulations of crystalline rubrene [23] and a discotic liquid crystal [32, 30] show that the most intensive Fourier component of the transfer-integral autocorrelation function corresponds to either underdamped [23] or overdamped [30] oscillations in the time domain. In other words, the spectral density (SD) of the carrier–phonon interaction, [33, 34] which is the only phonon-dependent quantity determining purely carrier dynamics under standard assumptions, [35] can be reasonably described within the Brownian-oscillator (BO) model. [36, 37, 33]
@@
When approached from the perspective of open quantum dynamics, [38, 39] phonon-limited charge transport is usually studied by tracking the relaxation of an initial nonequilibrium charge distribution towards equilibrium. [40, 41, 42, 43, 44] However, such approaches cannot access the time-dependent diffusion constant or frequency-dependent mobility, [45, 46] which are related to the finite-temperature time-dependent current autocorrelation function. [47, 48] Methods of open quantum dynamics have recently been used to compute autocorrelation function of the purely electronic current operator within the Holstein model. [49, 50, 44, 45, 51, 46] Analogous computations for the Peierls model are difficult because one has to express the expectation values involving the phonon-assisted current [52, 53, 54, 49] in terms of purely electronic quantities remaining upon integrating phonons out.
@@
This problem can be solved within the DEOM framework, [55, 56, 57, 58, 59] which we employ here to study carrier transport dynamics in a dissipative Peierls model. [21, 30] In contrast to most open quantum dynamics-based studies, [40, 42, 44] which employ the Drude–Lorentz (DL) SD to model the interaction of a carrier with low-frequency phonons, [33] we assume the more physically plausible [60] BO SD. We concentrate on model parameters representative of room-temperature hole transport along the maximum-conductivity direction in single crystals of rubrene. [23, 3] We use the generalized Wick’s theorem [55, 56, 57, 58, 59] to handle the phonon-assisted current and circumvent potential long-time numerical instabilities [61] by applying an appropriate DEOM closing scheme, [62, 24] which enables us to reliably access the diffusive dynamics and thus the low-frequency dynamical mobility.
@@
For weakly damped phonons, subdiffusive carrier dynamics remain limited to intermediate timescales, similarly to our HEOM results at zero damping. [25] The diffusive transport is then approached from the superdiffusive side, giving rise to a dynamical-mobility enhancement below the phonon frequency. As the damping is increased, the diffusive transport sets in from the subdiffusive side, and DEOM results are qualitatively similar to TLS predictions and experimental results, which feature a steady dynamical-mobility suppression below the DDP. We find that this change in the character of the long-time dynamics and low-frequency optical response occur in the region of underdamped BO, at dampings such that the phonon frequency remains well defined. Physically reasonable variations in the damping constant weakly affect the dc mobility, which remains within experimental bounds. Our results provide strong evidence that the long-time superdiffusive dynamics and low-frequency dynamical-mobility enhancements reported in minimal models [25, 21] are artifacts of the assumed undamped phonon dynamics.
@@
The paper is organized as follows. Section II introduces the model and the DEOM formalism. Section III discusses some technical aspects necessary to obtain reliable DEOM results, which we present in Sec. IV. We summarize our main conclusions in Sec. V.
@@ II.2 Spectral density of carrier–phonon interaction
which, in the limit of zero damping $\gamma_{0}$, reduces to the delta-like SD studied in Refs. 1, 3, 21, 24, 25. Microscopically, the damping can be traced back to the coupling of the undamped vibration to an Ohmic bath, [36, 37] which can be considered to stem from the quasicontinuum of neglected low-frequency vibrational modes that typically span the $(0-200)\>\mathrm{cm}^{-1}$ range. [30, 5] The BO SD can thus be regarded as an effective SD modulating transitions between electronic states that is equivalent to the Ohmic SD modulating transitions between vibronic (mixed vibrational and electronic) states. [66, 67, 68] As the SD conveniently combines information on the carrier–phonon interaction strength and phonon density of states, see Eq. (5), the BO SD can be considered as a convenient model of the influence of a continuous (as opposed to the delta-like [1, 3, 21, 24, 25]) density of phonon states on carrier dynamics. The quantity
@@
Figure 1: Frequency profile of (a) the BO SD [Eq. (6)] and (b) the BO friction spectrum $\mathcal{J}_{\mathrm{BO}}(\omega)/\omega$ for different values of the damping $\gamma_{0}$. The curves for $\gamma_{0}/\omega_{0}=0.05$ are scaled down by a factor of 4 for visual clarity. The inset in (a) compares the BO SD for $\gamma_{0}/\omega_{0}=4/\sqrt{17}$ to the DL SD [Eq. (9)] with $\gamma_{\mathrm{DL}}=\gamma_{0}/2$.
The BO is underdamped for $\gamma_{0}<\omega_{0}$, critically damped for $\gamma_{0}=\omega_{0}$, and overdamped for $\gamma_{0}>\omega_{0}$. The changes in the BO SD and the so-called frictional spectrum $\mathcal{J}_{\mathrm{BO}}(\omega)/\omega$ [69] with $\gamma_{0}$ are summarized in Figs. 1(a) and  1(b), respectively. The frictional spectrum, which essentially determines the free-bath propagator at high temperatures [Eq. (11) with $\beta\omega\ll 1$ whenever $\mathcal{J}_{\mathrm{BO}}(\omega)$ is appreciable], [31] peaks at a non-zero (zero) frequency for $\gamma_{0}<\omega_{0}/\sqrt{2}$ ($\gamma_{0}\geq\omega_{0}/\sqrt{2}$). [70]
@@
where $\gamma_{\mathrm{DL}}$ is the characteristic bath frequency. In contrast to the high-frequency behavior $\mathcal{J}_{\mathrm{BO}}(\omega)\propto\omega^{-3}$ of the BO SD, the DL SD exhibits a long high-frequency tail $\mathcal{J}_{\mathrm{DL}}(\omega)\propto\omega^{-1}$. Therefore, even in the high-temperature regime $2\pi T\gg\gamma_{\mathrm{DL}}$, to which the typically used values of $\gamma_{\mathrm{DL}}=40\>\mathrm{cm}^{-1}$ and $T=300\>\mathrm{K}=210\>\mathrm{cm}^{-1}$ belong, [40, 42, 44] the long tail of the DL SD implies that HEOM/DEOM computations have to take into account a number of Matsubara terms to achieve full convergence. [60] To reduce that number, Ishizaki [60] proposed that the DL SD in Eq. (9) be replaced with the BO SD in Eq. (6) with $\gamma_{0}=2\gamma_{\mathrm{DL}}$ and $\omega_{0}^{2}=\gamma_{0}^{2}+\epsilon^{2}$ with $|\epsilon|\lesssim\gamma_{0}/4$. The inset of Fig. 1(a) compares the DL SD to the corresponding BO SD for $|\epsilon|=\gamma_{0}/4$, so that $\gamma_{0}/\omega_{0}=4/\sqrt{17}$. In addition to lowering the number of Matsubara terms, Ishizaki’s approximation to DL SD circumvents unphysical short-time dynamics of mixed carrier–phonon quantities, which we discuss in Sec. IV.4.
@@ II.3 Dissipaton decomposition
with $\eta_{q_{2}q_{1}m}=\delta_{q_{1}\overline{q_{2}}}c_{m}/N$. Equation (12) somewhat differs from the standard DEOM prescription, which features $\delta_{m_{1}m_{2}}$ instead of $\delta_{m_{1}\overline{m_{2}}}$, see, e.g., Eq. (3.2) of Ref. 57 or Eq. (8) of Ref. 59. Our choice in Eq. (12) is motivated by the requirement that the dissipative theory smoothly reduces to the dissipationless case as $\gamma_{0}\to 0$. Then, we have demonstrated [24] that $f_{qm}$ are proportional to the creation and annihilation operators $b_{q}$ and $b_{\overline{q}}^{\dagger}$ of a single vibrational quantum. The same continuity argument implies that $f_{qm}^{\dagger}=f_{\overline{q}\>\overline{m}}$, while the results of Ref. 57 suggest that $f_{qm}^{\dagger}=f_{\overline{q}m}$. We emphasize that these subtle differences between our and original DEOM formulations exist only in the underdamped regime, [71] in which $\mu_{m}$s appear in complex-conjugated pairs. While we expect that the differences would affect dissipaton-resolved quantities, the quantities depending on $f_{qm}$ only through their sum $B_{q}$ [see, e.g., Eq. (16)] do not depend on whether Eq. (12) features $\delta_{m_{1}\overline{m_{2}}}$ or $\delta_{m_{1}m_{2}}$. Most of the previous DEOM-based studies dealt with the (strongly) overdamped regime, [72, 57, 58, 73] in which all $\mu_{m}$s are real ($\overline{m}=m$), so that Eq. (12) reduces to the standard DEOM prescription. The DEOM construction also needs the time-reversal counterpart of Eq. (12), which reads
@@ II.4 DEOM-based framework for transport properties
In Eq. (20), $V^{\times}O=[V,O]$, $V^{>}O=VO$, $V^{<}O=OV$, while $\mu_{\mathbf{n}}=\sum_{qm}n_{qm}\mu_{m}$. At each instant $t$, we combine Eqs. (14)–(16) and (17)–(19) to obtain

### v2

@@ Abstract
Optical conductivity in molecular semiconductors is suppressed in the terahertz region, featuring the displaced Drude peak that reflects carriers’ transient localization (TL) by slow intermolecular vibrations. Meanwhile, recent computations in minimal models evidence optical-conductivity enhancements below the characteristic vibrational frequency, which cannot be captured by the TL phenomenology. These models assume that the carrier’s hopping amplitude is modulated by a single undamped vibration. The modulation is, however, by many low-frequency modes, whose net effect can be approximated using a few effective damped oscillators. Here, we employ the dissipaton equations of motion (DEOM) method to compute the finite-temperature real-time current autocorrelation function in a one-dimensional model with Brownian-oscillator spectral density of nonlocal carrier–phonon interaction. We exploit the dissipaton algebra to handle the phonon-assisted current, reduce the method’s computational requirements by working in momentum space, and confirm that numerically stable transport dynamics are virtually independent of a specific DEOM closing scheme. With increasing damping, we find that DEOM optical-conductivity profiles become increasingly qualitatively similar to TL predictions. For parameters representative of room-temperature hole transport in single-crystal rubrene, we conclude that the TL phenomenology is established already in the underdamped-oscillator regime. Reasonable variations in the damping constant weakly affect the carrier mobility, which remains within experimental bounds. Overall, our results strongly suggest that optical-conductivity enhancements at very low frequencies are artifacts of the assumed delta-like phonon spectrum and rationalize the success of the TL phenomenology in describing experimental data.
@@ I Introduction
Recent quantum–classical [21] as well as our fully quantum numerically “exact” [24, 25] computations in the above-described one-dimensional Peierls (or Su–Schrieffer–Heeger) [26, 27, 28] model have suggested that TLS predictions are qualitatively inaccurate on long timescales (at low frequencies). For model parameters appropriate for room-temperature hole transport along the maximum-conductivity direction in single crystals of rubrene, [23]the authors of Refs. 25 and 21 evidence optical-response enhancements at very low frequencies. Meanwhile, most experimental low-frequency signatures align with TLS (Drude–Anderson) predictions, [9, 10, 11, 12] displaying a steady suppression of the optical response below the DDP, though there are exceptions. [8]
@@
The aforementioned minimal-model studies [25, 21] assume that the carrier hopping amplitude is modulated by a single phonon mode (also known as the “killer” mode [7, 29]) of frequency $\omega_{0}\approx 50\>\mathrm{cm}^{-1}$. Although the “killer” mode does have the dominant contribution to the total dynamic disorder, a more realistic model should acknowledge the presence of other low-frequency modes, typically quasi-continuously distributed over the $(0-200)\>\mathrm{cm}^{-1}$ range. [30] A viable strategy to tackle such a model is to assume that only a few (or even a single) effective oscillators are actually coupled to the carrier and to incorporate the effect of the remaining modes by subjecting the effective oscillators’ dynamics to a random force and the corresponding (via the fluctuation–dissipation relation) friction. [30, 31] Indeed, it was shown that the spectrum of the transfer-integral autocorrelation function emerging from molecular-dynamics simulations of a discotic liquid crystal [32, 30] can be accurately represented as a sum of three effective oscillators. While the most dominant oscillator for the material studied in Ref. 30 exhibits overdamped dynamics, one can argue [31, 21] that the most relevant oscillator for molecular crystals performs underdamped oscillations. [23] In other words, the spectral density (SD) of the carrier–phonon interaction, [33, 34] which is the only phonon-dependent quantity determining purely carrier dynamics under standard assumptions, [35] can be reasonably described within the Brownian-oscillator (BO) model. [36, 37, 33]
@@
When approached from the perspective of open quantum dynamics, [38, 39] phonon-limited charge transport is usually studied by tracking the relaxation of an initial nonequilibrium charge distribution toward equilibrium. [40, 41, 42, 43, 44] However, such approaches cannot access the time-dependent diffusion constant or frequency-dependent mobility, [45, 46] which are related to the finite-temperature time-dependent current autocorrelation function. [47, 48] Methods of open quantum dynamics have recently been used to compute the autocorrelation function of the purely electronic current operator within the Holstein model. [49, 50, 44, 45, 51, 46] Analogous computations for the Peierls model are difficult because one has to express the expectation values involving the phonon-assisted current [52, 53, 54, 49] in terms of purely electronic quantities remaining upon integrating phonons out. Focusing on models with a finite number of undamped phonon modes, [1, 3, 21] which lack explicit dissipation, we have recently succeeded in doing so using the hierarchical equations of motion (HEOM) method. [24, 25] The crux of our solution is the explicit expression of HEOM auxiliaries in terms of phonon creation and annihilation operators. [24] In models with a continuous distribution of phonon frequencies, [40, 42, 44] which are referred to as dissipative, such an expression is not yet available, to the best of our knowledge. [55, 56]
@@
Here, we rely on the DEOM formalism [57, 58, 59, 60, 61] to obtain numerically “exact” real-time quantum results in a dissipative Peierls model, [21, 30, 31] which one can use to assess the TLS and quantum–classical approximations. [21] While the dynamical equations of the DEOM formalism are identical to those of the HEOM formalism, the former provides the generalized Wick’s theorem, [57, 58, 59, 60, 61] which we use to handle the phonon-assisted current. We concentrate on model parameters representative of room-temperature hole transport along the maximum-conductivity direction in single crystals of rubrene, [23, 3]presenting one of the first applications of the generalized Wick’s theorem in an extended model system. In contrast to most open quantum dynamics-based studies, [40, 42, 44] which employ the Drude–Lorentz (DL) SD to model the interaction of a carrier with low-frequency phonons, [33] we assume the more formally appropriate [62] BO SD, whose underdamped variant is also more physically plausible for molecular crystals. [31] We circumvent potential long-time numerical instabilities [63] by applying an appropriate DEOM closing scheme, [64, 24] which enables us to reliably access the diffusive dynamics and thus the low-frequency dynamical mobility.
@@
For weakly damped phonons (a narrow spectrum of relevant modes), subdiffusive carrier dynamics remain limited to intermediate timescales, similarly to our HEOM [25] and quantum–classical [21] results at zero damping. The diffusive transport is then approached from the superdiffusive side, giving rise to a dynamical-mobility enhancement below the phonon frequency. As the damping is increased (the spectrum of relevant modes widens), the diffusive transport sets in from the subdiffusive side, and DEOM results are qualitatively similar to TLS predictions and experimental results. We find that this change in the character of the long-time dynamics and low-frequency optical response occurs in the region of underdamped BO, at dampings such that the central frequency of the relevant-phonon spectrum remains well defined. This rationalizes the success of the TLS in reproducing experimental low-frequency dynamical-mobility signatures in organic crystals, in which relevant phonon modes are narrowly distributed around the “killer” mode. [23, 31] Physically reasonable variations in the damping constant (the width of the relevant-phonon spectrum) weakly affect the dc mobility, which remains within experimental bounds and whose value can be reproduced reasonably well with the usual choice of the free parameter of the TLS. [65, 66] We relate this parameter to the damping parameter of the BO SD and find that the overall quality of TLS predictions for dynamical mobility improves with increasing damping. Then, the BO SD becomes increasingly similar to the DL SD, [67] and the need to treat the phonon-assisted current permits us to expose formal deficiencies of the DL SD, [62] which are not apparent in the related Holstein model, [41, 45, 46, 49, 50, 51] and showcase an application of the procedure to circumvent them. [62] Overall, our fully quantum results provide strong evidence that the long-time superdiffusive dynamics and low-frequency dynamical-mobility enhancements reported in minimal models [25, 21] are artifacts of the assumed delta-like phonon spectrum (undamped phonon dynamics).
@@
This paper is organized as follows. Section II introduces the model and the DEOM formalism. Section III discusses some technical aspects necessary to obtain reliable DEOM results, which we present in Sec. IV. We summarize our main conclusions in Sec. V.
@@ II.2 Spectral density of carrier–phonon interaction
which, in the limit of zero damping $\gamma_{0}$, reduces to the delta-like SD studied in Refs. 1, 3, 21, 24, 25. The quantity
@@
As the SD conveniently combines information on the carrier–phonon interaction strength and phonon density of states, see Eq. (5), the BO SD can be considered as a convenient model of the influence of a continuous (as opposed to the delta-like [1, 3, 21, 24, 25]) density of phonon states on carrier dynamics. Additionally, the BO model was argued to be relevant for one-dimensional organic semiconductors because it smoothly connects limiting transport regimes in the field of purely dynamic ($\gamma_{0}/\omega_{0}\to 0$) and purely static ($\gamma_{0}/\omega_{0}\to\infty$) disorder. [31, 71] The changes in the BO SD and the so-called frictional spectrum $\mathcal{J}_{\mathrm{BO}}(\omega)/\omega$ [72] with $\gamma_{0}$ are summarized in Figs. 1(a) and  1(b), respectively. The frictional spectrum, which essentially determines the free-bath propagator at high temperatures [Eq. (11) with $\beta\omega\ll 1$ whenever $\mathcal{J}_{\mathrm{BO}}(\omega)$ is appreciable], [31] peaks at a non-zero (zero) frequency for $\gamma_{0}<\omega_{0}/\sqrt{2}$ ($\gamma_{0}\geq\omega_{0}/\sqrt{2}$). [67] The underdamped regime $\gamma_{0}/\omega_{0}<1$, in which the width of the relevant spectrum (proportional to $\gamma_{0}$) is such that the vibrational frequency remains well defined, is relevant for organic crystals. [23, 31] The overdamped regime $\gamma_{0}/\omega_{0}>1$ is relevant for more statically disordered systems. [31, 30]
Figure 1: Frequency profile of (a) the BO SD [Eq. (6)] and (b) the BO frictional spectrum $\mathcal{J}_{\mathrm{BO}}(\omega)/\omega$ for different values of the damping $\gamma_{0}$. The curves for $\gamma_{0}/\omega_{0}=0.05$ are scaled down by a factor of 4 for visual clarity. The inset in (a) compares the BO SD for $\gamma_{0}/\omega_{0}=4/\sqrt{17}$ to the DL SD [Eq. (9)] with $\gamma_{\mathrm{DL}}=\gamma_{0}/2$.
@@
where $\gamma_{\mathrm{DL}}$ is the characteristic bath frequency. In contrast to the high-frequency behavior $\mathcal{J}_{\mathrm{BO}}(\omega)\propto\omega^{-3}$ of the BO SD, the DL SD exhibits a long high-frequency tail $\mathcal{J}_{\mathrm{DL}}(\omega)\propto\omega^{-1}$. Therefore, even in the high-temperature regime $2\pi T\gg\gamma_{\mathrm{DL}}$, to which the typically used values of $\gamma_{\mathrm{DL}}=40\>\mathrm{cm}^{-1}$ and $T=300\>\mathrm{K}=210\>\mathrm{cm}^{-1}$ belong, [40, 42, 44] the long tail of the DL SD implies that HEOM/DEOM computations have to take into account a number of Matsubara terms to achieve full convergence. [62] To reduce that number, Ishizaki [62] proposed that the DL SD in Eq. (9) be replaced with the BO SD in Eq. (6) with $\gamma_{0}=2\gamma_{\mathrm{DL}}$ and $\omega_{0}^{2}=\gamma_{0}^{2}+\epsilon^{2}$, with $|\epsilon|\lesssim\gamma_{0}/4$. The inset of Fig. 1(a) compares the DL SD to the corresponding BO SD for $|\epsilon|=\gamma_{0}/4$, so that $\gamma_{0}/\omega_{0}=4/\sqrt{17}$. In addition to lowering the number of Matsubara terms, Ishizaki’s approximation to DL SD circumvents unphysical short-time dynamics of mixed carrier–phonon quantities, which we discuss in Sec. IV.4.
@@ II.3 Dissipaton decomposition
with $\eta_{q_{2}q_{1}m}=\delta_{q_{1}\overline{q_{2}}}c_{m}/N$. Equation (12) somewhat differs from the standard DEOM prescription, which features $\delta_{m_{1}m_{2}}$ instead of $\delta_{m_{1}\overline{m_{2}}}$ [see, e.g., Eq. (3.2) of Ref. 59 or Eq. (8) of Ref. 61]. Our choice in Eq. (12) is motivated by the requirement that the dissipative theory smoothly reduces to the dissipationless case as $\gamma_{0}\to 0$. Then, we have demonstrated [24] that $f_{qm}$ are proportional to the creation and annihilation operators $b_{q}$ and $b_{\overline{q}}^{\dagger}$ of a single vibrational quantum. The same continuity argument implies that $f_{qm}^{\dagger}=f_{\overline{q}\>\overline{m}}$, while the results of Ref. 59 suggest that $f_{qm}^{\dagger}=f_{\overline{q}m}$. We emphasize that these subtle differences between our and original DEOM formulations exist only in the underdamped regime, [55] in which $\mu_{m}$s appear in complex-conjugated pairs. While we expect that the differences would affect dissipaton-resolved quantities, the quantities depending on $f_{qm}$ only through their sum $B_{q}$ [see, e.g., Eq. (16)] do not depend on whether Eq. (12) features $\delta_{m_{1}\overline{m_{2}}}$ or $\delta_{m_{1}m_{2}}$. Most of the previous DEOM-based studies dealt with the (strongly) overdamped regime, [73, 59, 60, 74] in which all $\mu_{m}$s are real ($\overline{m}=m$), so that Eq. (12) reduces to the standard DEOM prescription. The DEOM construction also needs the time-reversal counterpart of Eq. (12), which reads
@@ II.4 DEOM-based framework for transport properties
In Eq. (20), $V^{\times}O=[V,O]$, $V^{>}O=VO$, $V^{<}O=OV$, while $\mu_{\mathbf{n}}=\sum_{qm}n_{qm}\mu_{m}$. At each instant $t$, we combine Eqs. (14)–(19) to obtain